OpenStudy (babyslapmafro):

Please help me solve the following problem. Find the center and radius of the sphere, given the following equation. (Click to see)

OpenStudy (babyslapmafro):

\[2x^2-2x+2y^2-3y+2z^2+5z=2\]

OpenStudy (babyslapmafro):

I've solved problems like this before but I do not know what to do when there is a coefficient greater than one in front of the squared variable...

OpenStudy (anonymous):

dividing everything by two might make it easier to look at. but you're basically just completing the square for ever variable. you're also allowed to do, for 2x^2 - 2x: \[= (\sqrt{2}x - \frac{ 1 }{ \sqrt{2} })^2 - \frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

so the answer is pi :D

OpenStudy (anonymous):

let me know if you have questions. and for the location of the center, it's the same concept as 2d. for (x - h)^2, h is the center's x-component because x = h makes (x - h)^2 = 0 so our x component is when x = 1/2